# Regularization Through Reasoning: Systematic Improvements in Language Model Classification via Explanation-Enhanced Fine-Tuning

### v1

@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across diverse natural language processing tasks (Zhang et al., 2023), excelling particularly in text generation and complex reasoning scenarios (Mula et al., 2023). However, adapting these models for domain-specific classification tasks through fine-tuning presents unique challenges (Bai et al., 2022). A critical concern is ensuring that fine-tuned models develop robust decision-making processes rather than exploiting superficial correlations between inputs and labels (Aghajanyan et al., 2023). This challenge becomes particularly pronounced in conversational quality evaluation, where models must demonstrate nuanced understanding of multiple communication aspects simultaneously.
Building on recent advances in explanation-based learning (Lampinen et al., 2022; Zhao et al., 2024), this paper investigates how incorporating explanations during fine-tuning affects model performance on conversational quality assessment tasks. We focus on three fundamental dimensions of conversation quality: naturalness of language, comprehensiveness of responses, and adherence to topic. These dimensions represent critical components of effective human-computer interaction and are applicable across diverse conversational contexts (Zheng et al., 2023). For each dimension, models learn to predict ratings on a 5-point scale, with higher scores indicating superior quality.
Our experimental approach centers on fine-tuning a 7B parameter student model using training data generated from an ensemble of larger teacher models. This methodology leverages the knowledge of more capable teacher models while maintaining the computational efficiency of smaller student models for practical deployment (Ho et al., 2022). We systematically compare two fine-tuning strategies: training exclusively on class labels versus training on both class labels and natural language explanations that justify those labels. To ensure comprehensive evaluation, we test these approaches across 6 diverse conversational datasets, evaluating performance on 3 quality dimensions each, yielding 18 distinct dataset-task combinations that span various conversational contexts and domains.
@@ 2.1 Related Work
Recent advances in explanation-based fine-tuning have shown promising results across various domains. Aghajanyan et al. (2023) demonstrated that explanation-based fine-tuning makes models more robust to spurious correlations, leading to better generalization on challenging inference tasks. Similarly, Zhao et al. (2024) investigated the role of explanations in fine-tuning and found that explanation-augmented training consistently improves model performance compared to standard approaches.
The development of reasoning capabilities in smaller language models has been a key focus of recent research. Mula et al. (2023) showed that smaller models can achieve remarkable performance when trained with step-by-step explanations from larger teacher models, while Ho et al. (2022) demonstrated effective methods for teaching reasoning to small language models through explanation-based fine-tuning. These works highlight the potential for explanation-based approaches to bridge the performance gap between large and small models.
Lampinen et al. (2022) provided a formal framework for understanding when and how models can learn from explanations, identifying key factors that determine explanation effectiveness. Their theoretical analysis complements empirical findings and provides guidelines for designing explanation-based learning systems. The broader field of instruction tuning (Zhang et al., 2023) has also emphasized the importance of providing models with rich training signals beyond simple input-output pairs.
@@ 3 Experimental Setup
  • Movies: Cornell Movie Dialogs Corpus (Danescu-Niculescu-Mizil & Lee, 2011)
@@ 5.4 LoRA Weight Differences
Table 5 presents the normalized differences in LoRA weights between models fine-tuned with and without explanations, using the Frobenius norm. Our analysis of attention weight patterns and hidden representations (Danilevsky et al., 2023) reveals that the largest differences occur in the Query projection of attention layers, followed by the MLP Gate projection and Up projections. This suggests that explanations during fine-tuning primarily impact the model’s ability to attend to relevant input tokens and modulate the flow of information through the network’s gating mechanisms. The MLP layers, responsible for non-linear transformations, also exhibit substantial differences, particularly in the Gate projections that control the information flow. The Key and Value projections, responsible for encoding the input representations, exhibit smaller but non-negligible differences, indicating that explanations also influence how the model perceives and encodes input data. The Output projection, which combines attention outputs, shows relatively minor differences, implying that final output generation is less affected by explanation presence during fine-tuning.

### v2

# arXiv:2511.02044v2[cs.LG] 01 Mar 2026

@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across diverse natural language processing tasks (Zhang et al., 2023), excelling particularly in text generation and complex reasoning scenarios (Mitra et al., 2023). However, adapting these models for domain-specific classification tasks through fine-tuning presents unique challenges (Bai et al., 2022). A critical concern is ensuring that fine-tuned models develop robust decision-making processes rather than exploiting superficial correlations between inputs and labels (Ludan et al., 2023). This challenge becomes particularly pronounced in conversational quality evaluation, where models must demonstrate nuanced understanding of multiple communication aspects simultaneously.
Building on recent advances in explanation-based learning (Hase and Bansal, 2022; Ballout et al., 2024), this paper investigates how incorporating explanations during fine-tuning affects model performance on conversational quality assessment tasks. We focus on three fundamental dimensions of conversation quality: naturalness of language, comprehensiveness of responses, and adherence to topic. These dimensions represent critical components of effective human-computer interaction and are applicable across diverse conversational contexts (Zheng et al., 2023). For each dimension, models learn to predict ratings on a 5-point scale, with higher scores indicating superior quality.
Our experimental approach centers on fine-tuning a 7B parameter student model using training data generated from an ensemble of larger teacher models. This methodology leverages the knowledge of more capable teacher models while maintaining the computational efficiency of smaller student models for practical deployment (Magister et al., 2022). We systematically compare two fine-tuning strategies: training exclusively on class labels versus training on both class labels and natural language explanations that justify those labels. To ensure comprehensive evaluation, we test these approaches across 6 diverse conversational datasets, evaluating performance on 3 quality dimensions each, yielding 18 distinct dataset-task combinations that span various conversational contexts and domains.
@@ 2.1 Related Work
Recent advances in explanation-based fine-tuning have shown promising results across various domains. Ludan et al. (2023) demonstrated that explanation-based fine-tuning makes models more robust to spurious correlations, leading to better generalization on challenging inference tasks. Similarly, Ballout et al. (2024) investigated the role of explanations in fine-tuning and found that explanation-augmented training consistently improves model performance compared to standard approaches.
The development of reasoning capabilities in smaller language models has been a key focus of recent research. Mitra et al. (2023) showed that smaller models can achieve remarkable performance when trained with step-by-step explanations from larger teacher models, while Magister et al. (2022) demonstrated effective methods for teaching reasoning to small language models through explanation-based fine-tuning. These works highlight the potential for explanation-based approaches to bridge the performance gap between large and small models.
Hase and Bansal (2022) provided a formal framework for understanding when and how models can learn from explanations, identifying key factors that determine explanation effectiveness. Their theoretical analysis complements empirical findings and provides guidelines for designing explanation-based learning systems. The broader field of instruction tuning (Zhang et al., 2023) has also emphasized the importance of providing models with rich training signals beyond simple input-output pairs.
@@ 3 Experimental Setup
  • Movies: Cornell Movie Dialogs Corpus (Danescu-Niculescu-Mizil and Lee, 2011)
@@ 5.4 LoRA Weight Differences
Table 5 presents the normalized differences in LoRA weights between models fine-tuned with and without explanations, using the Frobenius norm. Our analysis of attention weight patterns and hidden representations (Ballout et al., 2023) reveals that the largest differences occur in the Query projection of attention layers, followed by the MLP Gate projection and Up projections. This suggests that explanations during fine-tuning primarily impact the model’s ability to attend to relevant input tokens and modulate the flow of information through the network’s gating mechanisms. The MLP layers, responsible for non-linear transformations, also exhibit substantial differences, particularly in the Gate projections that control the information flow. The Key and Value projections, responsible for encoding the input representations, exhibit smaller but non-negligible differences, indicating that explanations also influence how the model perceives and encodes input data. The Output projection, which combines attention outputs, shows relatively minor differences, implying that final output generation is less affected by explanation presence during fine-tuning.